Pete's Boats has beginning long-term debt of $840 and ending long-term debt of $790. The beginning and ending total debt balances are $1,220 and $1,360, respectively. The interest paid is $30. What is the amount of the cash flow to creditors? Multiple Choice −$110 $80 $110 −$80 $20
Answer #1

Net new long term debt=Ending long term debt-Beginning long term debt

=790-840=($50)

Hence cash flow to creditors=Interest paid-Net new long term debt

=30-(50)

which is equal to

=$80.